OpenStudy (anonymous):

You are setting up a part-time business with an initial investment of $15,000. The unit cost of the product is $11.80, and the selling price is $19.30. (a) Find equations for the total cost C and total revenue R for x units. (b) Find the break-even point by finding the point of intersection of the cost and revenue equations. (c) How many units would yield a profit of $1000?

OpenStudy (anonymous):

hi. I am confuse about some other question.please help me on this too.

OpenStudy (anonymous):

what about the others:D & please tell me the solution.

OpenStudy (anonymous):

I think it is correct.

OpenStudy (anonymous):

it is correct.tell me about the solution and the other.

OpenStudy (anonymous):

see i know the formula for cost is\[Total cost = Total Fixed Cost \times Total variable Cost\] \[Total Revenue = No. of Units Sold \times Price Per Unit\]

OpenStudy (anonymous):

second question is that when Total Cost will be equal to the revenue

OpenStudy (anonymous):

\[Break Even Point = \frac{Fixed Cost}{Selling Price - Variable Cost per Unit}\]